OpenStudy (anonymous):

express as a polynomial. (1 sec for equation)

OpenStudy (anonymous):

\[(6x^2yz^3-xy^2z)/(xyz)\]

OpenStudy (chaise):

\[\frac{6x^2yz^3−xy^2z}{xyz}\] = xyz(6xz^2-y)/xyz = 6xz^2-y Sorry, my equation editor has been all funny.

OpenStudy (anonymous):

How do you get the fraction the way you did? also thanks for the help

OpenStudy (chaise):

Write this into equation editor: frac{top}{bottom}

OpenStudy (anonymous):

Very nice i have not been able to figure that out for the longest